# TOGLL: Correct and Strong Test Oracle Generation with LLMs

### v1

@@ III-A RQ1: Selecting LLMs and prompts for oracle generation
In this research question, we fine-tune various large language models, as outlined in Section II-A3, utilizing six unique prompt formats discussed in Section II-A2. The purpose of this study is to determine the most effective combination of model and prompt format for the test oracle generation task. This approach allows us to understand the impact of prompt engineering on model efficacy and guides the selection of the most effective model-prompt pair for a deeper investigation.
@@ III-A1 Experimental Setup
For this study, we utilized the SF110 training dataset discussed in Section II-A1, which includes 159,073 input samples. Each sample contains a unique identifier, the method under test (MUT), a test prefix, and a docstring, from which we constructed six prompts as outlined in Section II-A2. For the fine-tuning process of the code LLMs, we employed a GPU setup consisting of 4 A100 GPUs, each with 40GB of memory. To facilitate parallel processing and accommodate larger models such as CodeGen-2B and PolyCoder-2.7B within the 40GB memory limit, we utilized PyTorch Accelerator and DeepSpeed [40]. The models were fine-tuned using a learning rate of 2.5e-5 and the AdamW optimizer. Each LLM was fine-tuned across all six prompts, and accuracy was computed on the validation dataset using the ‘exact match’ metric. This metric may underestimate the overall accuracy, as a generated oracle may not be an exact match but can still be correct, as illustrated in  4.
@@ III-A2 Results
Table I displays the specifics of each prompt—detailing the information used in each prompt—on the right side. P1 utilizes only the test prefix as input, prompting the model to generate either an assertion or an exception oracle. P2 adds a docstring in addition to the test prefix, while P3 includes a method signature. P4 includes both docstring and method signature after the test prefix. P5 includes the entire MUT code with the test prefix, whereas, P6 contains both the docstring and entire MUT code with the test prefix. On the left side of Table I, we present the performance of seven fine-tuned code models in generating test oracles for all six prompt types.
@@
Among the seven code models evaluated, CodeGen-350M emerged as the top performer, closely followed by CodeParrot-110M, both highlighted in bold. In our RQ2 analysis for unseen inference, we identified these models as leading candidate LLMs.
@@ III-B RQ2 : Assessing test oracles correctness
RQ1 investigates various large language models’ test oracle generation performance when they are fine-tuned on the training portion of the SF110 dataset and their accuracy is evaluated on the validation portion of the same dataset. However, a fine-tuned model may not generalize when applied to unseen data, a common challenge in LLM applicability [17]. To address this concern, we designed this study to generate test oracles for 25 unseen large-scale Java projects discussed in Section  II-B1. A test oracle is correct if it captures the expected behavior of a software system. Generated test oracles that do not compile, that disagree with software behavior, or that are vacuous are considered incorrect. We study the rate of success in generating correct oracles for both TOGLL and TOGA.
@@ III-B1 Experimental Setup
Test Oracle Generation. We generate test oracles for the unseen dataset with TOGLL, which is a fine-tuned CodeParrot-110M model operating on P6, which includes the test prefix, docstrings (if available), and the MUT. The unseen dataset consist of a total of 202,434 assertion prefixes requiring TOGLL to generate an assertion oracle for each input sample and 21,123 exception prefixes requiring TOGLL to generate an exception oracle. We have used a single A100 GPU with 10GB of memory to perform the inference task. The inference dataset keeps a record of detailed metadata so that the generated test oracles can be mapped back to the test class and the test prefix, enabling test validation to be performed later.
Oracle Correctness Validation. Unlike RQ1, which assessed accuracy using the ‘exact match’ metric, RQ2 involves executing the integrated test cases—combining the test prefix with the generated test oracles—to collect detailed metrics for each type of oracle. For both assertion and exception oracles, we compute success rates across all 25 artifacts using Equation  1, and compare these rates to those achieved by TOGA.
@@ III-B2 Results
In Figure 3, we present the performance of correct test oracle generation by our LLM-based method, TOGLL, and the deep-learning-based method, TOGA. The Y-axis displays success rate as a percentage. The X-axis enumerates the names of 25 artifacts. Each artifact is represented by two sets of vertical bars: blue for TOGLL and red for TOGA. Solid blue bars show TOGLL’s success rate in generating assertion oracles, while blue bars with diagonal lines illustrate its success in generating exception oracles. Similarly, solid red bars represent TOGA’s assertion oracle success rates, and red bars with diagonal lines show its performance in generating exception oracles.
Figure 3 shows that TOGLL achieved significantly higher success rate across all projects for both assertion and exception oracle generation. For example, for apache commons-number, TOGA could achieve a success rate of only 1%, whereas, TOGLL achieve significantly higher accuracy of 74.7%. Similarly, for http-request, TOGA achieved 14% where as TOGLL achieved 83% success rate. Statistical analysis reveals a significant difference in the average success rates, with TOGLL achieving a mean of 63% compared to TOGA’s mean of 16.4%. A t-test yielded a t-statistic of 12 and a p-value of 1.1e-11, indicating a statistically significant difference in performance between the two methods, with TOGLL outperforming TOGA in terms of correct test oracle generation.
For exception oracle, the performance different is even higher. TOGA’s average success rate is 18.86%, while LLM achieves a significantly higher average of 93.4%. A t-test further supports this distinction with a t-statistic of 24.8 and a p-value of 1.3e-18, highlighting TOGLL’s superior performance in generating exception oracles compared to TOGA.
@@ III-C RQ3: Investigating diversity in oracle generation
In RQ2, we present the overall performance of TOGLL in generating correct test oracles and compare it against SOTA deep-learning method. RQ3 delves into understanding the diversity of the generated test assertions. More specifically, we investigate whether during unseen inference TOGLL follows the assertion distribution observed during training and also to what extent the generated assertions exactly match the ground truth assertions. These metrics can provide insights into the diversity of the generated assertions. Diversity is an important aspect of test oracle generation, allowing TOGLL to complement developer written oracles or those generated by other tools and thereby enable detection of faults that would otherwise be undetected. In RQ4, we investigate the ability of TOGLL to detect unique faults.
@@ III-C1 Experimental Setup
To set up this experimental study, we utilize the training dataset and results from RQ2. We divided the assertions into seven categories, consisting of the most frequently used JUnit assertions investigated in previous studies [51, 10, 16]. We count the total number of assertions in each category and calculate the distribution relative to the total number of assertions in the training data. We follow a similar process to compute the categorical distribution of TOGLL generated assertions during unseen inference. Furthermore, we compute the percentage of TOGLL generated assertions that are an exact match for ground truth assertions in our dataset.
@@ III-C2 Results
There is a clear evidence that the distribution of oracles generated by TOGLL is quite different from the training data. The training data has a sharp peak for assertEquals, whereas the inference data has lower peaks for two categories. Moreover the tails of the distributions have a much higher share of samples indicating that generation of assertions is much less concentrated than training data. With regard to assertion category TOGLL is more diverse in generating oracle assertions.
We see that out of the 194k generated assertions, only 18k are exact matches, indicating that the generated assertion are diverse relative to provided assertions in the dataset. We observe this diversity can come multiple ways. For example, the generated assertion may come from a different category, like the first and third examples in Figure 4, or may vary the assertion arguments, like the second example in Figure 4. We observe that ground truth oracles do not check all variables within the prefix which gives TOGLL an opportunity to generate diverse assertion oracle. We see this in the first example where the TOGLL generated assertion is requiring that sort yield the same array, whereas the EvoSuite assertions just checks the length of the array. Similarly the second example shows a TOGLL generated assertion for the q1 variable, whereas EvoSuite’s assertion targets the q2 variable. For the third example, TOGLL generates a different type of oracle. Not only are these TOGLL generated assertions diverse, but they potential add bug detection power which we explore in the next RQ.
@@
RQ3 Finding: TOGLL can generate assertions that are diverse with respect to both the assertion statement used and the variables and expressions targeted for observation in those assertions.
III-D RQ4 : Assessing test oracle strength
TABLE III: Bug-Detection Performance of TOGLL vs. EvoSuite Generated Assertions.
@@ III-D RQ4 : Assessing test oracle strength
Correct test oracles are essential for aligning with program requirements, yet correctness alone may not suffice to detect diverse bugs. The strength of test oracles is equally crucial. In RQ4, we explore the strength of generated test assertions and comparing TOGLL to the SOTA EvoSuite and TOGA.
@@ III-D1 Experimental Setup
To set up this experiment, we generated three distinct test suites for each of the 25 unseen artifacts: TTOGLL, TES, TIO. TTOGLL comprises test cases with TOGLL-generated assertions, excluding all test cases with non-compiling and incorrect (i.e., false positive) assertions produced by TOGLL. We recorded the set of final test cases, including the test class and the unique test identifier. For constructing TES, we use the same set of tests, but the test assertions were generated by EvoSuite. For TIO, the same set of test cases was retained; however, this time, the test cases did not contain any test assertions. The three test suites share identical counts of test cases, test case prefixes, and assertion totals. The unique aspect distinguishing them is the method used for generating assertions.
We utilize mutation testing to introduce a wide array of bugs into the source code of 25 artifacts. Mutation testing is a widely used method for evaluating the effectiveness of test suites in detecting bugs. Unlike bug benchmarks such as Defects4J  [20] or QuixBugs [26], which are limited in both the number and diversity of faults, mutation testing enables thousands of diverse faults to be injected. Recent studies have also pointed out that Defects4J might inflate the perceived bug detection capabilities of test suites, further supporting the preference for mutation testing  [16]. We employ PIT [8], a state-of-the-art mutation testing tool for Java programs. PIT offers several advantages over other tools, as it generates meaningful mutants and fewer equivalent mutants [22, 23, 39]. We have used PIT-1.9.8 to generate a total of 69,793 strong mutants. For each mutated artifact, we run all three test suites to detect mutants. Mutants detected by TIO are those that can be detected through implicit assertions, meaning test prefixes alone suffice for their detection without the need for explicit assertions. Computing TIO allows our study to follow the recommendations of [16] by removing mutants killed by implicit oracles from study results. We calculate the total number of mutants detected by EvoSuite and TOGLL assertions as well as the unique mutants killed by oracles generated by each technique. Additionally, we also provide the total and unique mutants killed by TOGA.
@@ III-D2 Results
In Table III, the first Column shows the names of the 25 artifacts. The second and third Columns present the total and unique mutants detected by EvoSuite-generated assertions, respectively. The fourth and fifth columns present the total and unique mutants detected by TOGLL assertions, respectively. Row 27 shows the total mutants detected across all artifacts by each method, whereas row 28 presents the performance of TOGA, in detecting total and unique mutants.
@@
For all 25 artifacts, test cases were generated by EvoSuite, as discussed in Section II-B1. EvoSuite, a search-based dynamic approach, generates assertions based on executed program behavior and actual result values. Thus, it is expected to produce strong assertions capable of detecting a large number of mutants. TOGLL, in contrast, is a static approach that utilizes the test prefixes generated by EvoSuite, but lacks knowledge of any values computed during test execution. Despite this disadvantage, TOGLL-generated assertions detected 83% of the EvoSuite-detected mutants. This significantly outperformed TOGA, which detected only 43% of the EvoSuite-detected mutants under the same conditions. A t-test yielded a t-statistic of 4.12 and a p-value of 0.0003, indicating a statistically significant difference in performance between the TOGLL and TOGA in terms of total bug detection.
Figure 5: Examples illustrating that TOGLL generated assertions can detect unique mutants that EvoSuite generated assertions (ES) can not.
TOGLL generated assertions were able detect 1,023 mutants that EvoSuite missed, using the same set of input prefixes. This is a significant improvement over TOGA, which only identified 105 unique mutants. In total, TOGLL killed nearly twice the number of mutants as TOGA, but when viewed relative to TOGA as a baseline TOGLL killed 9.7 times the number of unique mutants. This demonstrates that LLM-based techniques can generate a significant numbers of strong test assertions.
We performed a qualitative study of the TOGLL assertions that could kill unique mutants. This study was opportunistic in that we look into the 1023 unique mutants, their corresponding mutated MUTs from smallest to largest and stopped when we found a handful of examples that illustrated oracle diversity. Figure 5 shows two examples from the joda-time and Apache commons-angle package. In the first example, test327 calls the calculatePrintedLength method which was mutated by ‘replacing long division with multiplication’, TOGLL checks the return variable int0 and was able to kill the mutant. EvoSuite checks a value that is independent of the execution of the MUT and has no chance to kill the mutant. The second example, TOGLL check that angle_Rad1 should not be null and thus was able to detect the mutant which replaced the return value with null. Here again the EvoSuite generated assertion checks a value that is independent of the MUT. While anecdotal these examples illustrate how TOGLL leverages the EvoSuite generated test prefix and produces valuable test oracles that EvoSuite can not. We plan to conduct a more systematic study of the qualitative characteristics of both test oracles that are unique to EvoSuite and unique to TOGLL to help understand their strengths and guide technique refinement.
@@
III-E Threats to Validity
In our large-scale study, we have used open-source and widely recognized dataset SF110 [13]. Yet, our results may not generalize across other datasets. To address this threat and ensure generalizability, we further evaluated the performance of our fine-tuned model on the unseen dataset from  [16].

### v2

@@ III-A RQ1: Selecting LLMs and Prompts for Oracle Generation
In this research question, we fine-tune various large language models, as outlined in Section II-A3, utilizing six unique prompts discussed in Section II-A2. The purpose of this study is to determine the most effective combination of model and prompt format for the test oracle generation task. This approach allows us to understand the impact of prompts on model efficacy and guides the selection of the most effective model-prompt pair for a deeper investigation.
@@ III-A1 Experimental Setup
For this study, we utilized the SF110 training dataset discussed in Section II-A1, which includes 159,073 input samples. Each sample contains a unique identifier, the method under test (MUT), a test prefix, and a docstring, from which we constructed six prompts as outlined in Section II-A2. For the fine-tuning process of the code LLMs, we employed a GPU setup consisting of 4 A100 GPUs, each with 40GB of memory. To facilitate parallel processing and accommodate larger models such as CodeGen-2B and PolyCoder-2.7B within the 40GB memory limit, we utilized PyTorch Accelerator and DeepSpeed [43]. The models were fine-tuned using a learning rate of 2.5e-5 and the AdamW optimizer. Each LLM was fine-tuned across all six prompts, and accuracy was computed on the validation dataset using the ‘exact match’ metric. This metric may underestimate the overall accuracy, as a generated oracle may not be an exact match but can still be correct, as illustrated in Figure  4.
@@ III-A2 Results
Table I displays the specifics of each prompt—detailing the information used in each prompt—on the right side. P1 utilizes only the test prefix as input, prompting the model to generate either an assertion or an exception oracle. P2 adds a docstring in addition to the test prefix, while P3 includes a method signature. P4 includes both docstring and method signature after the test prefix. P5 includes the entire MUT code with the test prefix, whereas, P6 contains both the docstring and entire MUT code with the test prefix. On the left side of Table I, we present the performance of seven fine-tuned code models in generating test oracles for all six prompts.
@@
Among the seven code models evaluated, CodeGen-350M emerged as the top performer, closely followed by CodeParrot-110M, both highlighted in bold. The three top performing prompts are P4, P5, P6. RQ2 is performed on the unseen OracleEval25 dataset. Our investigation shows that CodeParrot-110M generalizes better on the unseen data than CodeGen-350M. Also, although, P5 works better in this study we find that P6 works better on the OracleEval25 dataset, which has more Javadoc (for 75% samples) compared to the SF110 dataset (45%) and P6 can leverage those Javadoc to generate better test oracles.
@@ III-B RQ2: Assessing Test Oracles Correctness
RQ1 investigates various LLMs’ test oracle generation performance when they are fine-tuned on the training portion of the SF110 dataset and their accuracy is evaluated on the validation portion of the same dataset. However, a fine-tuned model may not generalize when applied to unseen data, a common challenge in LLM applicability [44]. To investigate the generalizability of our method TOGLL, we perform this study on the unseen OracleEval25 dataset discussed in Section  II-B1. A test oracle is correct if it captures the expected behavior of a software system. Generated test oracles that do not compile, that disagree with software behavior, or that are vacuous are considered incorrect. We study the rate of success in generating correct oracles for both TOGLL and TOGA.
@@ III-B1 Experimental Setup
We generate test oracles for the OracleEval25 unseen dataset using our method TOGLL, which is a fine-tuned CodeParrot-110M model operating on P6, which includes the MUT, test prefix and docstrings. This dataset consist of a total of 202,434 assertion prefixes requiring TOGLL to generate an assertion oracle for each input sample and 21,123 exception prefixes requiring TOGLL to generate an exception oracle. We have used a single A100 GPU with 10GB of memory to perform the inference task. The dataset keeps a record of the metadata so that the generated test oracles can be integrated with the test prefixes, enabling test validation to be performed later.
Unlike RQ1, which assessed accuracy using the ‘exact match’ metric, RQ2 involves executing the integrated test cases—combining the test prefix with the generated test oracles—to collect detailed metrics for each type of oracle. For both assertion and exception oracles, we compute success rates across all 25 projects using Equation  1, and compare these rates to those achieved by TOGA. Counting empty and non-compiling assertions is straightforward. For false positives, a failing test oracle can indicate either an incorrect oracle or a bug in the MUT. If the MUT is known to be correct, a failing test oracle signals a false alarm. We use this well-established technique for test oracle evaluation [5] to evaluate TOGLL-generated oracles. The OracleEval25 dataset includes the latest stable releases of projects with no known bugs. Therefore, any failure of a generated test oracle suggests it is incorrect.
@@ III-B2 Results
In Figure 3, we present and compare our LLM-based method, TOGLL’s correct test oracle generation performance with deep-learning-based method, TOGA. The Y-axis displays success rate as a percentage. The X-axis enumerates the names of 25 projects. Each project is represented by two sets of vertical bars: blue for TOGLL and red for TOGA. Solid blue bars show TOGLL’s success rate in generating assertion oracles, while blue bars with diagonal lines illustrate its success in generating exception oracles. Similarly, solid red bars represent TOGA’s assertion oracle success rates, and red bars with diagonal lines show its performance in generating exception oracles.
Figure 3 shows that TOGLL achieved significantly higher success rate across all projects for both assertion and exception oracle generation. For example, for apache commons-number, TOGA could achieve a success rate of only 1%, whereas, TOGLL achieve significantly higher accuracy of 74.7%. Similarly, for http-request, TOGA achieved 14% where as TOGLL achieved 83% success rate. Statistical analysis reveals a significant difference in the average success rates, with TOGLL achieving a mean of 63% compared to TOGA’s mean of 16.4%. A t-test yielded a t-statistic of 12 and a p-value of $1.1e^{-11}$, indicating a statistically significant difference in performance between the two methods, with TOGLL outperforming TOGA in terms of correct test oracle generation.
TOGA uses a non-ML method to generate five types of assertion candidates using predefined templates, and then an ML model ranks the candidates to find the best assertion. We found that around 62% of the time, the model could not generate any assertions. Additionally, TOGA uses the most frequent constants and variables to generate assertEquals oracles, resulting in high false positives. TOGLL is not restricted to certain types of assertions and does not use frequently appearing values to generate assertions. Moreover, TOGLL makes more effective use of Javadoc than TOGA, which showed almost no improvement with Javadoc. TOGLL’s training data and model backbone could also contribute to its superior performance, but this requires further investigation.
For exception oracle, the performance difference is even higher. TOGA’s average success rate is 18.86%, while TOGLL achieves a significantly higher average of 93.4%. A Mann-Whitney U test further confirms a statistically significant performance difference between TOGLL and TOGA, highlighting TOGLL’s superior performance in generating exception oracles compared to TOGA (U = 625.0, p = $1.4e^{-09}$). Javadoc typically specifies exception conditions for a MUT and TOGLL’s ability to leverage these Javadoc and the specified exception conditions is a key factor contributing to its superior performance.
@@ III-C RQ3: Investigating Diversity in Oracle Generation
Since TOGLL is fine-tuned on EvoSuite-generated test oracles, one concern is whether TOGLL simply learns to generate EvoSuite oracles. To this end, RQ3 investigates the diversity of the generated test assertions. More specifically, we investigate whether during unseen inference TOGLL follows the assertion distribution observed during training and also to what extent the generated assertions exactly match the ground truth assertions. These metrics can provide insights into the diversity of the generated assertions. Diversity is an important aspect of test oracle generation, allowing TOGLL to complement developer written oracles or those generated by other tools and thereby enable detection of faults that would otherwise be undetected. In RQ4 and RQ5, we investigate the ability of TOGLL to detect unique faults.
@@ III-C1 Experimental Setup
To set up this study, we utilize the training dataset and results from unseen inference in RQ2. We divided the assertions into seven categories, consisting of the most frequently used JUnit assertions investigated in previous studies [7, 17, 5]. We count the total number of assertions in each category and calculate the distribution relative to the total number of assertions in the training data. We follow a similar process to compute the categorical distribution of TOGLL generated assertions during unseen inference. Furthermore, we compute the percentage of TOGLL generated assertions that are an exact match for ground truth assertions in our dataset.
@@ III-C2 Results
There is a clear evidence that the distribution of assertions generated by TOGLL is quite different from the training data. The training data has a sharp peak for assertEquals, whereas the inference data has lower peaks for two categories. Moreover the tails of the distributions have a much higher share of samples indicating that generation of assertions is much less concentrated than training data. With regard to assertion category TOGLL is more diverse in generating assertion oracles.
We see that out of the 194k generated assertions, only 18k are exact matches, indicating that the generated assertion are diverse relative to provided assertions in the dataset. We observe this diversity can come multiple ways. For example, the generated assertion may come from a different category, like the first and third examples in Figure 4, or may vary the assertion arguments, like the second example in Figure 4. We observe that ground truth oracles do not check all variables within the prefix which gives TOGLL an opportunity to generate diverse assertion oracles. We see this in the first example where the TOGLL generated assertion is requiring that sort yield the same array, whereas the EvoSuite assertions just checks the length of the array. Similarly the second example shows a TOGLL generated assertion for the q1 variable, whereas EvoSuite’s assertion targets the q2 variable. For the third example, TOGLL generates a different type of oracle. Not only are these TOGLL-generated assertions diverse, but they potentially add unique bug detection power which we explore in the next RQs. We can conclude that fine-tuning with EvoSuite oracles does not constrain TOGLL to EvoSuite-like oracles, as it generates distinct yet correct assertions involving various assertion types, variables, and expressions, with only a 9.5% exact match rate. In RQ4, we demonstrate that these distinct oracles detect 1,023 unique bugs that EvoSuite could not detect (Table III). Pre-trained on a vast codebase, TOGLL’s LLM backbone has likely encountered both developer-written and automated assertions. By leveraging this pretraining knowledge, TOGLL can generate diverse yet accurate assertions with strong bug detection effectiveness.
@@
RQ3 Finding: TOGLL assertions are diverse with respect to both the assertion statement used and the variables and expressions targeted for observation in those assertions.
III-D RQ4: Assessing Mutant Detection Effectiveness
TABLE III: Mutant detection performance of TOGLL, EvoSuite and TOGA generated assertions.
@@ III-D RQ4: Assessing Mutant Detection Effectiveness
Correct test oracles are essential for aligning with program requirements; however, correctness alone may not suffice to detect a diverse range of bugs. The strength of test oracles is equally crucial. In RQ4, we investigate the mutant detection effectiveness of TOGLL-generated assertions and compare them with the SOTA EvoSuite and TOGA.
@@ III-D1 Experimental Setup
To set up this experiment, we generate three distinct test suites (TTOGLL, TES, TIO) for each of the 25 projects from the OracleEval25 dataset. TTOGLL comprises test cases with TOGLL-generated assertions, excluding all test cases with non-compiling and incorrect (i.e., false positive) assertions produced by TOGLL. We recorded the set of final test cases, including the test class and the unique test identifier. For constructing TES, we use the same set of tests, but the test assertions were generated by EvoSuite. For TIO, the same set of test cases was retained; however, this time, the test cases did not contain any test assertions. The three test suites share identical counts of test cases, test case prefixes, and assertion totals. The unique aspect distinguishing them is the method used for generating assertions.
We utilize mutation testing to introduce a wide array of bugs into the source code of 25 projects. Unlike bug benchmarks such as Defects4J  [36] or QuixBugs [45], which are limited in both the number and diversity of bugs, mutation test injects thousands of diverse bugs. We employ PIT [38], a state-of-the-art mutation testing tool for Java programs. PIT offers several advantages over other tools, as it generates meaningful mutants and fewer equivalent mutants [46, 47, 48]. We have used PIT-1.9.8 to generate a total of 69,793 strong mutants. For each mutated project, we run all three test suites to detect mutants. Mutants detected by TIO are those that can be detected through implicit assertions, meaning test prefixes alone suffice for their detection without the need for any explicit assertions. Computing TIO allows our study to follow the recommendations of [5] by removing mutants killed by implicit oracles from study results. We calculate the total number of mutants detected by EvoSuite, TOGLL and TOGA assertions as well as the unique mutants killed by assertions generated by each method.
@@ III-D2 Results
In Table III, Column 1 shows the names of 25 projects. Columns 2 and 3 present the total and unique mutants detected by EvoSuite-generated assertions, respectively. Columns 4 and 5 present the total and unique mutants detected by TOGLL assertions, respectively. Row 27 shows the total mutants detected across all projects by each method, whereas row 28 presents the performance of TOGA, in detecting total and unique mutants.
@@
For all 25 projects, test cases were generated by EvoSuite, as discussed in Section II-B1. EvoSuite, a search-based dynamic approach, generates assertions based on executed program behavior and actual result values. Thus, it is expected to produce strong assertions capable of detecting a large number of mutants. TOGLL, in contrast, is a static approach that utilizes the test prefixes generated by EvoSuite, but lacks knowledge of any values computed during test execution. Despite this disadvantage, TOGLL-generated assertions detected 83% of the EvoSuite-detected mutants. This significantly outperformed TOGA, which detected only 43% of the EvoSuite-detected mutants under the same conditions.
Figure 5: TOGLL generated assertions detecting unique mutants that EvoSuite generated assertions mcan not.
TOGLL generated assertions were able to detect 1,023 mutants that EvoSuite missed, using the same set of test prefixes. This is also a significant improvement over TOGA, which only identified 105 unique mutants. In total, TOGLL killed nearly twice the number of mutants as TOGA, but when viewed relative to TOGA as a baseline TOGLL killed 9.7 times the number of unique mutants. This demonstrates that TOGLL can generate a significant number of strong test assertions and highlights TOGLL’s capability to complement EvoSuite’s assertions.
We perform a qualitative study of the TOGLL assertions that could kill unique mutants. This study was opportunistic in that we look into the 1023 unique mutants, their corresponding mutated MUTs from smallest to largest and stopped when we found a handful of examples that illustrated oracle diversity. Figure 5 shows two examples from the joda-time and Apache commons-angle package. In the first example, test327 calls the calculatePrintedLength method which was mutated by ‘replacing long division with multiplication’, TOGLL checks the return variable int0 and is able to kill the mutant. EvoSuite checks a value that is independent of the execution of the MUT and has no chance to kill the mutant. The second example, TOGLL check that angle_Rad1 should not be null and thus was able to detect the mutant which replaces the return value with null. Here again the EvoSuite generated assertion checks a value that is independent of the MUT. While anecdotal these examples illustrate how TOGLL leverages the EvoSuite generated test prefixes and produces valuable test oracles that EvoSuite can not.
@@
III-E RQ5: Assessing Real Bug Detection Effectiveness
In RQ4, we evaluate TOGLL’s bug detection effectiveness using mutants. In this research question, we investigate TOGLL’s real bug detection effectiveness and compare its performance with TOGA
III-E1 Experimental Setup
To conduct this study, we utilize Defects4J[36], a benchmark dataset consisting of real Java bugs. To ensure a fair comparison with TOGA, we follow the exact same experimental setup: the same Docker container environment, the same input dataset, and the exact same set of scripts provided in the TOGA replication package[49]. The Defects4J dataset consists of 374 input samples, each with a MUT, a test prefix, and Javadoc documentation. Of these, 304 require an assertion and 70 require an exception oracle. For this dataset, we generate oracle predictions using our method, TOGLL. Next, TOGLL generated test oracles are integrated with the respective test prefixes. These integrated test cases (test prefix + oracle) are then executed on both the buggy and fixed program versions. A bug is considered detected if a test passes on the fixed version but fails on the buggy version, indicating that the test oracle correctly aligns with the expected behavior and is capable of detecting the buggy behavior. To understand the results better, we first present the oracle classification performance of both methods in Table IV, followed by their bug detection performance in Table V.
TABLE IV: Comparison of oracle classification performance between TOGLL and TOGA on the Defects4J dataset. Metrics include Precision (P), Recall (R), F1-Score (F1), and Accuracy (Accu.).
Approach Exception Assertion Overall
P R F1 P R F1 Accu. P R
TOGA .14 .12 .13 .83 .86 .84 .74 .49 .49
TOGLL 1 1 1 1 1 1 1 1 1
III-E2 Results
Table IV presents the oracle classification performance of both methods on the Defects4J dataset. In this classification task, each method determines which type of oracle is needed for a given input prefix. The table shows performance metrics for both classes: exception and assertion, as well as the overall metrics.
TOGA exhibits very poor precision, recall, and F1-score for the class of test prefixes requiring an exception oracle. In contrast, our method, TOGLL, demonstrates a very high classification accuracy. In RQ2, the exception oracle classification accuracy averaged 93.4% for 222k input samples. In this smaller study of 374 inputs it achieved an accuracy of 100%.
Regarding assertion oracles, although TOGA correctly classified that an assertion oracle is needed for 84% of the inputs, it generated an explicit assertion for only 40% of them. For 60% of inputs, TOGA could not generate any assertions. On the other hand, TOGLL correctly classified that an assertion oracle is needed with 100% accuracy for the Defects4J study, and for 97% of them, it generated an assertion oracle.
Table V shows the bug detection performance of both approaches. Column 2 shows the total number of bugs detected by exception oracles. TOGLL detected a total of 27 bugs with its exception oracles, whereas TOGA detected only 5. These findings align with the data in Table IV, which shows that TOGA’s exception oracle generation precision is only 0.14. These results are also consistent with the RQ2 findings, which show a significant performance difference between TOGA and TOGLL, indicating a similar trend across multiple datasets.
Column 3 shows the total number of bugs detected by explicit assertion oracles (non-empty assertions). For the input test prefixes requiring an assertion oracle, TOGA generated assertions for only 40% of the inputs (121 out of 304), detecting a total of 26 bugs. In contrast, TOGLL generated assertions for 97% of the inputs, detecting a total of 37 bugs.
Column 4 shows the number of bugs detected by test prefixes alone. TOGA could not generate assertions for 183 out of 304 prefixes. Simply running the test prefixes on the buggy versions detected 36 bugs – these are detected by the implicit oracles of the Java Runtime System that expects no uncaught exception. TOGLL did not generate an assertion for only 9 prefixes, with only 1 bug detected by running those prefixes.
As suggested by recent prior work [5, 50], bugs detected by implicit oracles when executing test prefixes throwing uncaught exceptions should not be considered as contributions of a test oracle generation method. We followed this guidance in RQ4 and applying it here by excluding bugs found in this way reveals that TOGLL detects 64 bugs, which is 106.5% more bugs than the 31 bugs detected by TOGA. Even if one were to count the bugs detected by implicit oracles, TOGLL detects 16% more bugs than TOGA on Defects4J.
TABLE V: Defects4J bug detection performance of TOGLL vs. TOGA. A bug can be detected by different type of oracles.
Approach Bug Detected By FP
Exception Oracle Explicit Assertion Oracle EvoSuite Test Prefix Only Total Unique
TOGA 5 26 36 56 0.45
TOGLL 27 37 1 65 0.42
RQ5 Finding: When excluding the bugs that are detected by test prefixes alone, TOGA detected a total of 31 Defects4J bugs with its exception and explicit assertion oracles. In comparison, TOGLL detected 64 bugs and improvement of 106%.
III-F Threats to Validity
In our large-scale study, we have utilized open-source and widely recognized SF110 dataset [25]. Yet, our results may not generalize across other datasets. To ensure generalizability, we further evaluated TOGLL on the unseen OracleEval25  [5] and Defects4J [36] dataset.